OpenStudy (anonymous):

find the area of the region inside both r = 7sin(theta) and 7sin(2theta) in the first quadrant

OpenStudy (anonymous):

You integrate with respect to theta.

OpenStudy (anonymous):

I think it is something like \[ \int _{\theta_1}^{\theta_2} \frac12\left[r(\theta)\right]^2~d\theta \]

OpenStudy (anonymous):

the limit of integration is what I have difficult with

OpenStudy (anonymous):

I think there will be two integrals

OpenStudy (anonymous):

You have to figure out which \(r\) is greater.

OpenStudy (anonymous):

\[ 7\sin(2\theta) = 14\sin(\theta)\cos(\theta) \]

OpenStudy (anonymous):

\[ 14\sin(\theta)\cos(\theta) - 7\sin(\theta) = 0\\ 7\sin(\theta)(2\cos(\theta)-1) = 0 \]

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok, I got one solution to be pi/3

OpenStudy (anonymous):

The bounds are going to be \(0\to\pi/2\) But at \(\pi/3\), we need to do a switch?

OpenStudy (anonymous):